# Understanding the Training and Generalization of Pretrained Transformer for Sequential Decision Making

### v1

@@ C.1 More experiments on training dynamics
Setup. In Figure 4 (a) and (b), we consider two tasks: a linear bandits task (with $2$-dimensional action space) and a newsvendor task (with $4$-dimensional contexts). Both tasks have infinite support of the prior environment distribution $\mathcal{P}_{\gamma}$ (i.e., infinite possible environments in both the pretraining and testing). The details of the tasks can be found in Appendix E.2. For each task, we independently train two transformer models by Algorithm 1. For both models, the training parameters are identical except for the $M_{0}$: the first curve (blue, thick line) uses $M_{0}=50$, while the second curve (orange, dashed line) uses $M_{0}=130$, meaning no transformer-generated data is utilized during training. All other parameters follow the configuration detailed in Appendix E.1. The figure shows the mean testing regret at each training iteration across 128 randomly sampled environments from $\mathcal{P}_{\gamma}$, with the shaded areas representing the standard deviation. We conduct the same experiment on the multi-armed bandits task (with infinite possible environments and $20$ arms) for a different architecture for sequential decision making, Decision Pretrained Transformer (DPT) [Lee et al., 2024], in Figure 4 (c). We refer to the original paper for details on the DPT’s architecture.
@@ C.2 More experiments on matchings of $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$
We plot the actions generated from $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$ on the same environment with the same contexts in Figure 5 across dynamic pricing and newsvendor tasks. Each subfigure has a different number of possible environments or possible demand types (which are shared in both the pretraining and testing phases). The population-level differences $\{\texttt{TF}_{\hat{\theta}}(h_{t})-\texttt{Alg}^{*}(h_{t})\}_{t=1}^{T}$ are presented in Figure 6. We observe that across different tasks: (1) $\texttt{TF}_{\hat{\theta}}$ nearly matches $\texttt{Alg}^{*}$, but (2) the matchings are not perfect. The actions from $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$ are not exactly the same, and their differences increase as the underlying problems become more complex, as seen by comparing the first and second rows in Figure 6. This latter observation numerically supports the conditions in Proposition 4.4.
Setup. For Figure 5, each subfigure is based on a sampled environment with a sampled sequence of contexts $\{X_{t}\}_{t=1}^{30}$ from the corresponding task. For tasks in (a),(c), the data generation process follows the description detailed in Appendix E.2, and the architecture of $\texttt{TF}_{\hat{\theta}}$ and the definition of $\texttt{Alg}^{*}$ can be found in Appendix E.1 and Appendix E.3 respectively. For tasks in (b), (d), there are 16 environments included in the support of $\mathcal{P}_{\gamma}$, where 8 environments have demand functions of the linear type and the other 8 have demand functions of the square type (see Appendix E.2 for the definitions of these two types). Both the pretraining and testing samples are drawn from $\mathcal{P}_{\gamma}$, i.e., from these 16 environments. The remaining setups for these tasks follow the methods provided in Appendix E.2.
@@
C.3 Comparison with oracle posterior algorithms
In this subsection, we study the behavior of $\texttt{TF}_{\hat{\theta}}$ by comparing regret performances on $\texttt{TF}_{\hat{\theta}}$ and the algorithms utilizing the posterior distribution $\mathcal{P}(\gamma|H)$ of environments. We name the later algorithms as the oracle posterior algorithms as they are assumed to know the posterior distribution $\mathcal{P}(\gamma|H)$ (or equally, the prior distribution $\mathcal{P}(\gamma)$).
We first present a general framework for these oracle posterior algorithms in Algorithm 2: posterior averaging, sampling, and median. For simplicity of notation, we omit the subscript $t$ and only consider a finite space $\Gamma$ of environments, while the algorithms can be easily extended to the infinite case. For the posterior median algorithm, we assume the action space $\mathcal{A}\subseteq\mathbb{R}$.
@@ C.3 Comparison with oracle posterior algorithms
Figure 7: Performances of $\texttt{TF}_{\hat{\theta}}$ and oracle posterior algorithms. The numbers in the legend bar are the final regret at $t=100$ and the shaded areas indicate the $90\%$ (empirical) confidence intervals.
Figure 7 presents the testing regret and action suboptimality for $\texttt{TF}_{\hat{\theta}}$ compared to oracle posterior algorithms. Generally, all algorithms exhibit good regret performance. However, we observe that $\texttt{TF}_{\hat{\theta}}$ is (i) superior to posterior sampling in multi-armed bandits, as shown in (a), and (ii) outperforms posterior averaging in linear bandits, as shown in (c).
@@
These observations illustrate that $\texttt{TF}_{\hat{\theta}}$ can outperform oracle posterior algorithms in certain scenarios. This further demonstrates that the advantage of $\texttt{TF}_{\hat{\theta}}$ over benchmark algorithms is not solely due to its use of prior knowledge about the task. Rather, $\texttt{TF}_{\hat{\theta}}$ discovers a new decision rule that achieves better short-term regret than oracle posterior algorithms. It can be more greedy than posterior sampling while exploring more than posterior averaging.
@@ C.4 $\texttt{TF}_{\hat{\theta}}$ as a solution to model misspecification
In contrast, $\texttt{TF}_{\hat{\theta}}$ offers a potential solution to model misspecifications. By generating pretraining samples from all possible types of environments (e.g., newsvendor models with both linear and non-linear demand functions), $\texttt{TF}_{\hat{\theta}}$ leverages its large capacity to make near-optimal decisions across different types. Figure 8 demonstrates the performance of $\texttt{TF}_{\hat{\theta}}$ trained on tasks with two types of demand functions: linear and square (definitions provided in Appendix E.2). It compares $\texttt{TF}_{\hat{\theta}}$ with benchmark algorithms designed solely for linear demand cases. The performance is tested in environments where demand functions (i) can be either type with equal probability (first column); (ii) are strictly square type (second column); and (iii) are strictly linear type (third column). The benchmark algorithms face model misspecification issues in the first two cases.
@@
Setup. For both the pricing and newsvendor tasks, during the pretraining phase of $\texttt{TF}_{\hat{\theta}}$, the pretraining data are generated from two types of demand functions: linear and square, with each type having a half probability. The support of $\mathcal{P}_{\gamma}$ is infinite, meaning the parameters in the demand functions are not restricted. The context dimension is 6 for the pricing task and 4 for the newsvendor task. The generating details and definitions of these two types of demands are provided in Appendix E.2. Additionally, when generating data using $\texttt{TF}_{\theta_{m}}$ in Algorithm 1, the sampled environment also has an equal probability of having a linear or square demand type.
Each subfigure in Figure 8 is based on 100 runs. We consider three different cases during testing: (i) the environment can have either the square or linear demand type (each with half probability); (ii) the environment can only have the linear demand type; and (iii) the environment can only have the square demand type. In the first case, the testing environment is sampled in the same way as the pretraining environment, while in the other two cases, we only include the sampled environments with the specified demand type until the total testing samples reach 100.
@@ C.5 Additional results on $\texttt{TF}_{\hat{\theta}}$ performances
This subsection presents further results on the performance of $\texttt{TF}_{\hat{\theta}}$. Specifically, we compare $\texttt{TF}_{\hat{\theta}}$ with benchmark algorithms on (i) simple tasks, which include only 4 possible environments in both pretraining and testing samples (Figure 9), and (ii) more complex tasks, which involve either 100 possible environments (Figure 10 (a), (b)) or 16 environments with two possible types of demand functions (Figure 10 (c), (d)).
Figures 9 and 10 illustrate the consistently and significantly superior performance of $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$ across all tasks compared to the benchmark algorithms. These results underscore the advantage of leveraging prior knowledge about the tested environments.
@@
For the tasks in Figure 10 (c), (d), there are 16 environments included in the support of $\mathcal{P}_{\gamma}$, where 8 environments have demand functions of the linear type and the other 8 have demand functions of the square type (see Appendix E.2 for the definitions of these two types). Both the pretraining and testing samples are drawn from $\mathcal{P}_{\gamma}$, i.e., from these 16 environments. The remaining setups for these tasks follow the methods provided in Appendix E.2.

### v2

@@ C.1 More experiments on training dynamics
C.1.1 Ablation study on $f$ and $\kappa$
In this section, we present ablation studies to explore the impact of two key factors: the decision function $f$ used for generating pre-training data, and the mix ratio $\kappa$ applied in Algorithm 1. Figure 7 summarizes the results, where we pre-train and test models in multi-armed bandits problems defined in Appendix E.2, and only tune $f$ or $\kappa$ during the pre-training and keep all other parameters the same as in Appendix E.1.2.
  • Effect of $f$: We evaluate the influence of different decision functions $f$ on the performance of the transformer in a multi-armed bandit task. Three types of decision functions are considered: the one we designed, as defined in Appendix E.1.2 , the UCB algorithm for multi-armed bandits [Lattimore and Szepesvári, 2020], and a 50/50 mixture of both. The training dynamics for each function are shown in Figure 7(a). We observe that all three approaches achieve similar final performance once training ends. While it is possible that some decision functions could lead to faster convergence or better performance after pre-training, we did not encounter any training failures due to the choice of $f$.
  • Effect of $\kappa$: We also conduct an experiment to numerically investigate the impact of different mix ratios $\kappa$. The results are presented in Figure 7(b). We test four values of $\kappa$ $(10\%,33\%,66\%,90\%)$, keeping all other hyperparameters constant for a multi-armed bandit task. The findings show that transformer’s performance remains robust across different values of $\kappa$, except when $\kappa=0.9$, where performance declines. This supports our earlier discussion in Algorithm 1, highlighting the importance of mixing a non-trivial amount of self-generated data (i.e., non-trivial value for $1-\kappa$).
(a) Impact of $f$
(b) Impact of $\kappa$
Figure 7: (a) The effect of $f$ (decision function for generating the pre-training data), where we mix the UCB algorithm [Lattimore and Szepesvári, 2020] and decision function designed in Appendix E.1.2 with different ratios to create the pre-training data. (b) The effect of $\kappa$ (ratio of samples generated by $f$ in the mixed training phase).
C.1.2 Additional results on Figure 2 (a)
@@
Setup. In Figure 8 (a) and (b), we consider two tasks: a linear bandits task (with $2$-dimensional action space) and a newsvendor task (with $4$-dimensional contexts). Both tasks have infinite support of the prior environment distribution $\mathcal{P}_{\gamma}$ (i.e., infinite possible environments in both the pre-training and testing). The details of the tasks can be found in Appendix E.2. For each task, we independently train two transformer models by Algorithm 1. For both models, the training parameters are identical except for the $M_{0}$: the first curve (blue, thick line) uses $M_{0}=50$, while the second curve (orange, dashed line) uses $M_{0}=130$, meaning no transformer-generated data is utilized during training. All other parameters follow the configuration detailed in Appendix E.1. The figure shows the mean testing regret at each training iteration across 128 randomly sampled environments from $\mathcal{P}_{\gamma}$, with the shaded areas representing the standard deviation. We conduct the same experiment on the multi-armed bandits task (with infinite possible environments and $20$ arms) for a different architecture for sequential decision making, Decision Pre-trained Transformer (DPT) [Lee et al., 2024], in Figure 8 (c). We refer to the original paper for details on the DPT’s architecture.
@@ C.2 More experiments on matchings of $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$
We plot the actions generated from $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$ on the same environment with the same contexts in Figure 9 across dynamic pricing and newsvendor tasks. Each subfigure has a different number of possible environments or possible demand types (which are shared in both the pre-training and testing phases). The population-level differences $\{\texttt{TF}_{\hat{\theta}}(H_{t})-\texttt{Alg}^{*}(H_{t})\}_{t=1}^{T}$ are presented in Figure 10. We observe that across different tasks: (1) $\texttt{TF}_{\hat{\theta}}$ nearly matches $\texttt{Alg}^{*}$, but (2) the matchings are not perfect. The actions from $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$ are not exactly the same, and their differences increase as the underlying problems become more complex, as seen by comparing the first and second rows in Figure 10. This latter observation numerically supports the conditions in Proposition 4.4.
Setup. For Figure 9, each subfigure is based on a sampled environment with a sampled sequence of contexts $\{X_{t}\}_{t=1}^{30}$ from the corresponding task. For tasks in (a),(c), the data generation process follows the description detailed in Appendix E.2, and the architecture of $\texttt{TF}_{\hat{\theta}}$ and the definition of $\texttt{Alg}^{*}$ can be found in Appendix E.1 and Appendix E.3 respectively. For tasks in (b), (d), there are 16 environments included in the support of $\mathcal{P}_{\gamma}$, where 8 environments have demand functions of the linear type and the other 8 have demand functions of the square type (see Appendix E.2 for the definitions of these two types). Both the pre-training and testing samples are drawn from $\mathcal{P}_{\gamma}$, i.e., from these 16 environments. The remaining setups for these tasks follow the methods provided in Appendix E.2.
@@
C.3 Regret comparison with $\texttt{Alg}^{*}$
In this subsection, we study the behavior of $\texttt{TF}_{\hat{\theta}}$ by comparing regret performances on $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$, which utilizes the posterior distribution $\mathcal{P}(\gamma|H)$ of environments. As shown in Section 4, $\texttt{Alg}^{*}$’s are defined as the Bayes-optimal decision functions.
Algorithm 2 presents a general framework for the Bayes-optimal decision functions/algorithms in Example 4.2: posterior averaging, sampling, and median. For simplicity of notation, we omit the subscript $t$ and only consider a finite space $\Gamma$ of environments, while the algorithms can be easily extended to the infinite case. For the posterior median algorithm, we assume the action space $\mathcal{A}\subseteq\mathbb{R}$.
@@ C.3 Regret comparison with $\texttt{Alg}^{*}$
Figure 11: Performances of $\texttt{TF}_{\hat{\theta}}$ and Bayes-optimal decision functions. The numbers in the legend bar are the final regret at $t=100$ and the shaded areas indicate the $90\%$ (empirical) confidence intervals.
Figure 11 presents the testing regret and action suboptimality for $\texttt{TF}_{\hat{\theta}}$ compared to the Bayes-optimal decision functions/algorithms. Generally, all algorithms exhibit good regret performance. However, we observe that $\texttt{TF}_{\hat{\theta}}$ is (i) superior to posterior sampling in multi-armed bandits, as shown in (a), and (ii) outperforms posterior averaging in linear bandits, as shown in (c).
@@
These observations illustrate that $\texttt{TF}_{\hat{\theta}}$ can outperform Bayes-optimal decision functions in certain scenarios. This further demonstrates that the advantage of $\texttt{TF}_{\hat{\theta}}$ over benchmark algorithms is not solely due to its use of prior knowledge about the task. Rather, $\texttt{TF}_{\hat{\theta}}$ discovers a new decision rule that achieves better short-term regret than oracle posterior algorithms. It can be more greedy than posterior sampling while exploring more than posterior averaging.
@@ C.4 $\texttt{TF}_{\hat{\theta}}$ as a solution to model misspecification
In contrast, $\texttt{TF}_{\hat{\theta}}$ offers a potential solution to model misspecifications. By generating pre-training samples from all possible types of environments (e.g., newsvendor models with both linear and non-linear demand functions), $\texttt{TF}_{\hat{\theta}}$ leverages its large capacity to make near-optimal decisions across different types. Figure 12 demonstrates the performance of $\texttt{TF}_{\hat{\theta}}$ trained on tasks with two types of demand functions: linear and square (definitions provided in Appendix E.2). It compares $\texttt{TF}_{\hat{\theta}}$ with benchmark algorithms designed solely for linear demand cases. The performance is tested in environments where demand functions (i) can be either type with equal probability (first column); (ii) are strictly square type (second column); and (iii) are strictly linear type (third column). The benchmark algorithms face model misspecification issues in the first two cases.
@@
Setup. For both the pricing and newsvendor tasks, during the pre-training phase of $\texttt{TF}_{\hat{\theta}}$, the pre-training data are generated from two types of demand functions: linear and square, with each type having a half probability. The support of $\mathcal{P}_{\gamma}$ is infinite, meaning the parameters in the demand functions are not restricted. The context dimension is 6 for the pricing task and 4 for the newsvendor task. The generating details and definitions of these two types of demands are provided in Appendix E.2. Additionally, when generating data using $\texttt{TF}_{\theta_{m}}$ in Algorithm 1, the sampled environment also has an equal probability of having a linear or square demand type.
Each subfigure in Figure 12 is based on 100 runs. We consider three different cases during testing: (i) the environment can have either the square or linear demand type (each with half probability); (ii) the environment can only have the linear demand type; and (iii) the environment can only have the square demand type. In the first case, the testing environment is sampled in the same way as the pre-training environment, while in the other two cases, we only include the sampled environments with the specified demand type until the total testing samples reach 100.
@@ C.5 Additional results on $\texttt{TF}_{\hat{\theta}}$ performances
This subsection presents further results on the performance of $\texttt{TF}_{\hat{\theta}}$. Specifically, we compare $\texttt{TF}_{\hat{\theta}}$ with benchmark algorithms on (i) simple tasks, which include only 4 possible environments in both pre-training and testing samples (Figure 13), and (ii) more complex tasks, which involve either 100 possible environments (Figure 14 (a), (b)) or 16 environments with two possible types of demand functions (Figure 14 (c), (d)).
Figures 13 and 14 illustrate the consistently and significantly superior performance of $\texttt{TF}_{\hat{\theta}}$ and $\texttt{Alg}^{*}$ across all tasks compared to the benchmark algorithms. These results underscore the advantage of leveraging prior knowledge about the tested environments.
@@
For the tasks in Figure 14 (c), (d), there are 16 environments included in the support of $\mathcal{P}_{\gamma}$, where 8 environments have demand functions of the linear type and the other 8 have demand functions of the square type (see Appendix E.2 for the definitions of these two types). Both the pre-training and testing samples are drawn from $\mathcal{P}_{\gamma}$, i.e., from these 16 environments. The remaining setups for these tasks follow the methods provided in Appendix E.2.
C.6 Performance on longer testing horizon
Our paradigm can be generalized to a longer testing horizon (beyond the length of the pre-training data) by introducing a context window. Specifically, we define a context window size $W$ and limit the input to the last $\min\{W,t\}$ timesteps of the input sequence $H_{t}$ to predict $a^{*}_{t}$ during both the pre-training and testing phases. By considering only the latest $\min\{W,t\}$ timesteps, we enable the model to handle testing sequences of any length, even when $t>T$. This technique is commonly used in the literature (e.g., Chen et al. [2021]).
In this part, we evaluate $\texttt{TF}_{\hat{\theta}}$’s generalization ability when the testing horizon is longer than that seen during pre-training. As for a shorter testing horizon, $\texttt{TF}_{\hat{\theta}}$ can simply run until the end of the horizon, and previous results demonstrate strong performances in such cases, particularly when the horizon is less than 100. Figure 15 shows the performance when the testing horizon is extended from 100 (pre-training horizon) to 200 in the newsvendor problem. The results demonstrate that $\texttt{TF}_{\hat{\theta}}$ generalizes well in this extended horizon scenario, even in the censored demand setting where exploration is essential. Its actions remain nearly optimal beyond $t=100$, resulting in lower regret compared to benchmark algorithms. For this experiment, all the pre-training and testing setups follow Appendix E, except the testing horizon length.
(a) Regret
(b) Action suboptimality
Figure 15: Performance under horizon generalization in newsvendor problems, where the pre-training samples have a horizon of 100 but the testing samples have a horizon of 200. It shows the average out-of-sample regret (first row) and action suboptimality, i.e., $|a^{*}_{t}-\texttt{Alg}(H_{t})|$, (second row) of $\texttt{TF}_{\hat{\theta}}$ against benchmark algorithms. The numbers in the legend bar are the final regret at $t=200$.
C.7 Out-of-distribution performance
In this section, we evaluate $\texttt{TF}_{\hat{\theta}}$’s generalization ability under out-of-distribution (OOD) conditions, using dynamic pricing problems as the test case.
(a) $\sigma^{2}=0.1$ ($<$ pre-training)
(b) $\sigma^{2}=0.2$ ($=$ pre-training)
(c) $\sigma^{2}=0.3$ ($>$ pre-training)
(d) $\sigma^{2}=0.1$ ($<$ pre-training)
(e) $\sigma^{2}=0.2$ ($=$ pre-training)
(f) $\sigma^{2}=0.3$ ($>$ pre-training)
Figure 16: OOD performance under different testing noise variances, which may deviate from the pre-training variance of $\sigma^{2}=0.2$. It shows the average out-of-sample regret (first row) and action suboptimality, i.e., $|a^{*}_{t}-\texttt{Alg}(H_{t})|$, (second row) of $\texttt{TF}_{\hat{\theta}}$ against benchmark algorithms. The numbers in the legend bar are the final regret at $t=100$.
We test $\texttt{TF}_{\hat{\theta}}$ in three different problem settings with varying testing noise variances while keeping the pre-training variance as $0.2$ for all of them: $\sigma^{2}=0.2$, which matches the noise variance used in pre-training; $\sigma^{2}=0.1$, which is smaller than that used in pre-training; and $\sigma^{2}=0.3$, which is larger. Figure 16 provides the results. We do not observe any significant sign of failure in $\texttt{TF}_{\hat{\theta}}$’s OOD performance: across all three settings, the benchmark algorithms consistently incur higher regret than $\texttt{TF}_{\hat{\theta}}$. Although $\texttt{TF}_{\hat{\theta}}$ shows some variation in mean regret across OOD settings (specifically, lower regret when $\sigma^{2}=0.2$ and higher regret when $\sigma^{2}=0.3$), this is partially due to the varying “difficulty” of the underlying tasks. As expected, higher noise variances, which need more data for accurately estimating the demand function compared to lower variance cases, lead to worse performance for all algorithms tested. However, these variations should not be interpreted as evidence of $\texttt{TF}_{\hat{\theta}}$ failing to handle OOD issues. In fact, all benchmark algorithms show similar performance fluctuations under these conditions, further demonstrating $\texttt{TF}_{\hat{\theta}}$’s good OOD performances. For this experiment, all the pre-training and testing setups follow Appendix E, except the noise variance in the testing samples.
C.8 Impact of network architecture
Here we investigate the effect of different network architectures on performance. Specifically, we replace the transformer/GPT-2 architecture used in $\texttt{TF}_{\hat{\theta}}$ with Long Short-Term Memory (LSTM) [Hochreiter, 1997].
We maintain the overall architecture of $\texttt{TF}_{\hat{\theta}}$, as described in Appendix E.1 except for replacing the transformer/GPT-2 module with LSTM. We evaluate two LSTM variants: a 5-layer LSTM and a 12-layer LSTM (for comparison, the tested transformer also has 12 layers). Typically, LSTM architectures should be shallower than transformers in practice, which is why we also include the 5-layer version. Other hyperparameters like embedding space dimension are kept the same across all models. We pre-train and test these models on dynamic pricing problems (see Appendix E.2 for the experimental setup), and the pre-training procedure is identical to the one outlined in Appendix 3.
Figure 17 presents the testing regret for the three models. The results indicate no significant difference between the 5-layer and 12-layer LSTM models. However, the transformer consistently performs better than both LSTM variants. This outcome aligns with the transformer’s superior performance in other domains, such as natural language processing, and suggests that transformers are more effective for sequential decision making tasks as well.
Figure 17: Comparison of the average out-of-sample regret between using the Transformer and LSTM architectures.
C.9 Ablation study on model size and task complexity
In this section, we explore the performance changes when tuning both model size and task complexity. Specifically, we adjust the number of layers to control the model size and the problem dimension (the dimension of the context) in a dynamic pricing task to control the task complexity, where the number of unknown parameters is twice the context dimension. Since the optimal reward may scale differently across dimensions, we provide a relatively fair comparison by evaluating the advantage of $\texttt{TF}_{\hat{\theta}}$ relative to the best-performing benchmark algorithm. This advantage is defined as the reward improvement rate of $\texttt{TF}_{\hat{\theta}}$ compared to the best benchmark algorithm (the one among ILSE, CILS, and TS that achieves the highest reward). Thus, a positive value indicates $\texttt{TF}_{\hat{\theta}}$ outperforms all benchmark algorithms, and a larger value indicates a better performance of $\texttt{TF}_{\hat{\theta}}$. We test three model sizes (4, 8, and 12 layers) and three problem dimensions (4, 10, and 20), while keeping other hyperparameters the same. The results are presented in Figure 18.
Figure 18: The advantage of $\texttt{TF}_{\hat{\theta}}$ compared to the best benchmark algorithm in dynamic pricing, across different model size (number of layers) and problem complexity (problem dimensions).
From Figure 18, we observe the following: (i) $\texttt{TF}_{\hat{\theta}}$ consistently outperforms the benchmark algorithms, and this advantage grows as the complexity of the problem increases. (ii) Increasing model size consistently improves the performance of $\texttt{TF}_{\hat{\theta}}$ across all problem dimensions, indicating that larger models are generally preferred. This highlights $\texttt{TF}_{\hat{\theta}}$’s superior ability to handle more complex tasks and these observations align with the scaling law [Kaplan et al., 2020] for large language models.